b) Find the cylindrical coordinates of the points where the Cartesian coordinates are (2)

i) )8,6,6(
ii) ( 1,1,2 )
1
Expert's answer
2020-05-21T12:42:26-0400

Correspondence between cylindrical "(\rho,\phi,z)" and Cartesian ("x,y,z)" is

"\rho=\sqrt{x^2+y^2},\phi=\arctan( y/x), z=z" then

i) "\rho=\sqrt{8^2+6^2}=10, \phi=\arctan(6/8)\approx 37^0,z=6"

ii)"\rho=\sqrt{1^2+1^2}=\sqrt{2},\phi=\arctan1=45^0,z=2"